OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3,-4) PLEASE HELP ITS ESSENTIAL THAT I PASS

OpenStudy (amistre64):

what line equation forms do you ahve to work with?

OpenStudy (anonymous):

i have to put it in point slope form and i only have the info that i listed

OpenStudy (amistre64):

you need to know a few different forms of line equations or make a generic slope intercept to work with

OpenStudy (anonymous):

ok ill try

OpenStudy (amistre64):

point slope form is what you need eh; do you recall the general equation for that?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

hmm, well given a point (a,b) and a slope m, the general set up is: y-b = m(x-a)

OpenStudy (anonymous):

ok but i really dont understand this so

OpenStudy (amistre64):

are you given a point and a slope?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then when given a point (a,b) and a slope m, the general set up is: y-b = m(x-a) just use the numbers you have and fill in the abm parts

OpenStudy (anonymous):

so m is the slope and y is the point

OpenStudy (amistre64):

m is the slope yes, and the point itself defines the values for a and b; (a,b)

OpenStudy (anonymous):

ok thx

OpenStudy (amistre64):

y-b = m(x-a) has a slope of -3: m= -3 y-b = -3(x-a) and passes through the point (3,-4); so what are the values for (a,b)?

OpenStudy (anonymous):

so it would be y- -4=-3(x-3) correct?

OpenStudy (amistre64):

correct, and the -- is just a +

OpenStudy (anonymous):

ok i appreciate it thx

OpenStudy (amistre64):

good luck :)